OpenStudy (anonymous):

when 1.4613 g of an organic compound containing Fe, C, H , O was burned in 02(oxygen) . 2.7333 g of CO2 AND 0.78222 g of H20 were produced. in a separate experiment to determine the mass of iron 0.4676 g, the compound yielded 0.1056 g of Fe2O3. what is the empirical formula of the compuond?

OpenStudy (koikkara):

\(Hint:\) \(\small\rm\color{green}{Check~out~this~similar~question~from~yahoo~!!}\) When 1.7358g of an organic iron compound containing Fe, C, H, and O was burned in O2, 3.2467g of CO2 and 0.92916g of H2O were produced. \(\small\rm\color{blue}{Answer!}\) (0.06996 g Fe2O3) / (159.6887 g Fe2O3/mol) x (2 mol Fe / 1 mol Fe2O3) x (55.8452 g Fe/mol) / (0.3096 g) = 0.158049 = 15.8049% Fe (3.2467g CO2) / (44.00964 g CO2/mol) x (1 mol C / 1 mol CO2) x (12.01078 g C/mol) = 0.886065 g C (0.92916 g H2O) / (18.01532 g H2O/mol) x (2 mol H / 1 mol H2O) x (1.007947 g H/mol) = 0.103972 g H 15.8049% of 1.7358 g = 0.274341 g Fe Use the Law of Conservation of Mass: (1.7358 g total) - (0.274341 g Fe) - (0.103972 g H) - (0.886065 g C) = 0.471422 g O Convert to moles: (0.274341 g Fe) / (55.8452 g Fe/mol) = 0.00491253 mol Fe (0.886065 g C) / (12.01078 g C/mol) = 0.0737725 mol C (0.103972 g H) / (1.007947 g H/mol) = 0.103152 mol H (0.471422 g O) / (15.99943 g O/mol) = 0.0294649 mol O Divide by the smallest number of moles: (0.00491253 mol Fe) / 0.00491253 mol = 1.0000 (0.0737725 mol C) / 0.00491253 mol = 15.017 (0.103152 mol H) / 0.00491253 mol = 20.998 (0.0294649 mol O) / 0.00491253 mol = 5.9979 Round to the nearest whole numbers to find the empirical formula: FeC15H21O6 \(\small\rm\color{blue}{This~is~how~you~solve~this~problem~!}\) \(But,\) \(\Large\rm\color{green}{Now~you~can~try~it~out~with~your~question!}\)

OpenStudy (koikkara):

@YamadaTasnim \(Hope~It~helps~!!\)

OpenStudy (koikkara):

\(Hint: Complete~Question~\) When 1.7358g of an organic iron compound containing Fe, C, H, and O was burned in O2, 3.2467g of CO2 and 0.92916g of H2O were produced. In a separate experiment to determine the mass percent of iron , 0.3096g of the compound yielded 0.06996g of Fe2O3. What is the empirical formula of the compound.